OpenStudy (anonymous):

HELP PLEASE URGENT Determine if the graph is symmetric about the x-axis, the y-axis, or the origin. r = 4 cos 5θ

OpenStudy (anonymous):

Hint! The cosine graph is an even Function

OpenStudy (anonymous):

so x axis?

OpenStudy (anonymous):

YES!

OpenStudy (anonymous):

Thank you! DO YOU HAVE TIME FOR ANYMORE?

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

Find the center, vertices, and foci of the ellipse with equation 2x2 + 8y2 = 16.

OpenStudy (anonymous):

Ok the equation of an ellipse has either form (x+h)^2/a^2+(y+k)^2/b^2=1 note the result it's always 1! so before you do anything check it's in this form and tell me what your new equation should look like...

OpenStudy (anonymous):

i have no idea how to do that like none of this makes any sense haha

OpenStudy (anonymous):

ok you know how the equation of an ellipse should look like right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

it should look like

OpenStudy (anonymous):

I want you to look it carefully and you should always remember it!

OpenStudy (anonymous):

okay i have seen that before

OpenStudy (anonymous):

see the 1 after the = sign that's important

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

to make your equation 2x2 + 8y2 = 16 in the form we've seen the first step is to divide(both sides) by 16 do you see why

OpenStudy (anonymous):

that would make our equation x2/8 + y2/2=1 which is in the form we've just seen